The equation y=-1777x+27153 can be used to predict the number y of gun deaths in the United States x years after 2000, that is, x=0 corresponds to 2000, x=3 correspnds to 2003, x=5 corresponds to 2005, and so on. Predict the number of gun deaths in 2005 and 2009. In what year will the gun deaths be 7606?
Answer by jorel1380(3719)

-1777(5)+27153=-8885+27153=18268;
-1777(9)+27153=-15993+27153=11160;
----------------------------
7606=27153-1777x
1777x=19547
x=11
Gun deaths will be 7606 in 2011..